1. Th e supply of loanable funds: comes from households that consume all of their income results from the desire to accumulate wealth for retirement or for major future expenditures c. is inversely related to the interest rate d. does not depend on the interest rate 2. Both consumer demand and investment demand for loanable funds will be: directly related to the interest rate inversely related to the interest rate c. unrelated to the interest rate A decrease in expected rates of return would lead to: a. an increase in the demand for loanable funds b. c. d. an increase in the quantity of borrowing a decrease in the equilibrium interest rate create an equilibrium in the bond market 4. Generally, interest rates will be lower on loans: a. that have high risk b. that have a longer term c. that have a smaller relative cost to make d. that is issued retail institutions 5. If the nominal interest rate is 8% and the rate of inflation is 3%, the real interest rate is: a, 11% b.7% с. 5% d.3% 6. The present value of future income: a. is the discounted value in the future b. is less than the total of the future income stream c. is greater than the total of the future income stream d. is equal to the total of the future income stream 7. The present value of $50,000 to be received three years from today, assuming a 3% interest rate is: a. $54,500 b. $50,000 c. $45,757 d. $44,872 . The present value of $50,000 to be received three years from today, assuming a 7% interest rate is: a. $40,815 b. $41,322
Answer #1

1..)

Supply of loanable fund:

  • Saving
  • Disinvestments
  • Dishoarding.

Right answer: (B)

2)

Investment demand for loanable fund falls when interest rate rises. Interest rate is a kind of cost for investors.

Hence, right answer: ( B)

3)

Decrease in expected rate of return suggests that now less investment shall be made. Hence, now demand for loanable fund also falls. Fall in demand will pull down interest rate.

Right answer: (C)

4)

Right answer: (C)

Reduce cost of lending would invariably lead to fall in interest rate as well. High risky loan are given at higher interest rate.

5)

Real interest rate = Nominal interest rate – inflation rate.

= 8 -3

= 5 %

Right answer ( C)

6)

Right answer: ( B)

When future income is discounted, it total value falls. Thus, present value of future income would be less than future income stream.

7)

Future income = P (1+r)^n

= 50,000 ( 1+0.03)^3

= 50,000*1.09

= 54,500

Right answer: (A)